A 20.0-L volume of methane gas is heated from 25°C to 70°C at constant pressure. What is the final volume of the gas?


1
Expert's answer
2022-05-17T04:15:04-0400

Given:

V1 = 20.0 L

T1 = 25°C + 273.15 = 298.15 K

P1 = P2 = constant

V2 = unknown

T2 = 70°C + 273.15 = 343.15 K


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final volume of the gas, solve the equation for V2:

V2 = V1T2 / T1

V2 = (20.0 L × 343.15 K) / (298.15 K) = 23.0186 L = 23.0 L

V2 = 23.0 L


Answer: The final volume of the gas is 23.0 L